### v1

# DeepWriter: A Fact-Grounded Multimodal Writing Assistant Based On Offline Knowledge Base

Abstract
Large Language Models (LLMs) have demonstrated remarkable capabilities in various applications. However, their use as writing assistants in specialized domains like finance, medicine, and law is often hampered by a lack of deep domain-specific knowledge and a tendency to hallucinate. Existing solutions, such as Retrieval-Augmented Generation (RAG), can suffer from inconsistency across multiple retrieval steps, while online search-based methods often degrade quality due to unreliable web content. To address these challenges, we introduce DeepWriter, a customizable, multimodal, long-form writing assistant that operates on a curated, offline knowledge base. DeepWriter leverages a novel pipeline that involves task decomposition, outline generation, multimodal retrieval, and section-by-section composition with reflection. By deeply mining information from a structured corpus and incorporating both textual and visual elements, DeepWriter generates coherent, factually grounded, and professional-grade documents. We also propose a hierarchical knowledge representation to enhance retrieval efficiency and accuracy. Our experiments on financial report generation demonstrate that DeepWriter produces high-quality, verifiable articles that surpasses existing baselines in factual accuracy and generated content quality.
DeepWriter: A Fact-Grounded Multimodal Writing Assistant Based On Offline Knowledge Base
Song Mao1, Lejun Cheng1,2, Pinlong Cai1, Guohang Yan1, Ding Wang1, Botian Shi1, 1Shanghai AI Lab, 2Peking University Correspondence: caipinlong@pjlab.org.cn
1 Introduction
The rapid advancement of Large Language Models (LLMs) and Multimodal Large Language Models (MLLMs) has unlocked a wide array of applications, from programming and education to complex task automation. Commercial models like Gemini DeepMind (2025), Claude 3.5 Anthropic (2024) and GPT-4o OpenAI (2025) have shown proficiency in leveraging web search and deep reasoning to assist users with real-world problems.
In this paper, we investigate the application of LLMs as sophisticated writing assistants. While the context windows of modern LLMs are sufficient for processing entire novels Team et al. (2024); MiniMax et al. (2025); Yang et al. (2025), their application in professional writing often remains at a superficial level of data processing and summarization Wu et al. (2025). In specialized domains such as finance, medicine, and law, LLMs frequently fail to generate expert-level responses, and are prone to factual hallucinations. Two primary paradigms have emerged to mitigate these issues. The first, Retrieval-Augmented Generation (RAG) Lewis et al. (2020), connects LLMs to external knowledge bases. However, RAG can struggle with maintaining coherence and consistency across long-form documents, as multi-turn retrieval may introduce disjointed information Gu et al. (2025). The second paradigm involves using online search to gather relevant information Chen et al. (2024); Sun et al. (2025), but the variable quality of web content can compromise the overall quality and reliability of the generated document.
Figure 1: The overall framework of DeepWriter.
To overcome these limitations, we propose DeepWriter, a multimodal, long-context writing assistant that operates on a curated, offline corpus. Unlike systems that primarily gather information from structured web pages, DeepWriter is engineered to deeply mine knowledge from information-rich but unstructured sources like PDFs. Our pipeline utilizes advanced document processing toolkits to extract and meticulously organize text, tables, and images with their corresponding metadata.
This foundation allows DeepWriter to structurally organize the retrieved information for multimodal generation. Where standard RAG systems struggle to place visual elements effectively, we introduce a novel multimodal reranking system to intelligently determine the most appropriate location for images and charts within the text. Furthermore, to ensure the final output is fully verifiable and trustworthy, DeepWriter excels at generating and citing multimodal content with fine-grained granularity. Instead of citing an entire source file, which is a common practice that makes validation intractable for large documents, our system tracks and attributes information at the paragraph level. This approach guarantees that the generated document is not only coherent and professional but also factually grounded and reliable, directly addressing the core challenge of hallucination in LLMs.
We test DeepWriter on financial report generation task. Experiment results show that DeepWriter, with a simple yet effective framework, can achieve comparable performance with other leading open-source long-context models.
The key contributions of our work include:
  1. 1.
    Hierarchical Knowledge Representation: We proposed a three-level (knowledge-chunk-document) representation that enables efficient multi-granularity retrieval while maintaining precise source attribution capabilities.
  2. 2.
    Structured Writing Pipeline: We developed a comprehensive pipeline that includes query rewriting, task decomposition, outline generation, iterative retrieval, section-by-section writing, and reflection mechanisms to ensure quality and coherence.
  3. 3.
    Multimodal Content Integration: We introduced an interleaved image-text generation approach with relevance scoring and placement optimization algorithms that seamlessly integrate visual and textual content.
  4. 4.
    Fine-grained Citation System: We implemented a three-level citation system (document, paragraph, sentence) that provides precise source attribution and enables comprehensive fact verification.
2 Related Work
Our work is situated at the intersection of three research areas: AI-powered research assistants, automated writing assistants, and Retrieval-Augmented Generation (RAG).
Research Assistants
Several projects have explored the integration of LLMs with public data sources for scientific question-answering and writing. Systems like OpenScholar Asai et al. (2024) and ScholarQA Singh et al. (2025) aim to assist researchers by surveying and synthesizing information. DeepResearcher trained a LLM-based deep research agents end-to-end through scaling reinforcement learning in real-word environments Zheng et al. (2025). DeepWriter extends this line of work by focusing on long-form, professional document generation from a controlled, offline corpus, with an emphasis on multimodal content.
Retrieval-Augmented Generation (RAG)
RAG is a powerful technique for grounding LLM outputs in external knowledge, addressing issues of outdated information and hallucination. Innovations like GraphRAG Edge et al. (2025), LightRAG Guo et al. (2024), and KAG Liang et al. (2024) have introduced more sophisticated retrieval mechanisms. However, generating long, coherent documents with RAG remains a challenge. DeepWriter’s structured pipeline and hierarchical knowledge representation are designed to improve coherence and factual consistency in long-form generation tasks.
Writing Assistants
The concept of an AI writing assistant is popular, with a focus on improving human productivity by generating text in various styles. Frameworks like OmniThink have explored a "slow-thinking" approach to generate articles based on search engine results Xi et al. (2025). However, OmniThink did not address multimodal information or user customization. Other related systems like STORM Shao et al. (2024) and CO-STORM Jiang et al. (2024) focus on generating outlines and then "filling in the blanks" but rely on web searches, which can introduce noise. RAPID can generate comprehensive and knowledge-intensive articles based on RAG Gu et al. (2025). QRAFT is an LLM-based agentic framework that mimics the writing workflow of human fact-checkers Sahnan et al. (2025).  Xiong et al. (2025) proposed a general agent framework that achieves human like adaptive writing through recursive task decomposition and dynamic integration of three fundamental task types, i.e. retrieval, reasoning, and composition. DeepWriter differentiates itself by using a curated knowledge base and integrating multimodal content generation.
3 Methodology
3.1 Task Definition
Given a collected domain-specific corpus $\mathcal{K}=\{D_{1},\dots,D_{m}\}$ and a user-provided query or topic $Q$, our goal is to generate a long document $P$ with the following characteristics:
  1. 1.
    Multimodal Content: The document should include relevant images, tables, and charts derived from the corpus.
  2. 2.
    Factual Grounding: All claims, figures, and critical viewpoints must be accompanied by citations pointing to their source in the corpus.
  3. 3.
    Semantic Coherence: The document must be logically consistent and flow naturally from one section to the next.
Compared with existing writing assistants, DeepWriter aims to thoroughly put information together. Since prompting for innovation may cause hallucination, in this paper, we require the agent to give less or simply not pose any novel conclusion.
3.2 System Architecture
The DeepWriter pipeline consists of an offline processing stage and an online generation stage, as illustrated in the Fig 1.
Offline Processing Stage
During the offline stage, we aim to preprocess corpus and store them in an easy-to-retrieve format in a knowledge base $\mathcal{B}$.
In the data processing stage, we aim to extract structured information from unstructured file formats such as PDF. We first use Fitz [^1] https://github.com/pymupdf/PyMuPDF or MinerU Wang et al. (2024) to extract text, tables, images from the original files. During the process, we keep all file and page metadata to avoid information loss when performing fact-checking. After extracting information from files, we choose appropriate chunk size to avoid chunks that are too small or too large. We also use advanced VLM such as Qwen2.5-VL Bai et al. (2025) to generate detailed image or table captions for multimodal retrieval.
We structure the extracted information into a hierarchical database. This involves a three-level representation (chunk-page-document) to facilitate efficient, multi-granularity retrieval. The top "document" level contains abstract concepts, such as year or domain information, the "page" level is an intermediate level, which trades offs the efficiency and granularity of retrieval, the final "chunk" level is composed of knowledge pieces, holding the raw source information. The data processing hierarchy is shown in Fig 2.
Figure 2: Hierarchical structure of data processing. The document is decomposed into pages, and each page is further divided into text chunks. This multi-level organization enables fine-grained and scalable data analysis.
Algorithm 1 DeepWriter Online Generation Stage Pipeline
0:  User query $Q$, Offline knowledge base $\mathcal{K}$
0:  Generated document $P$ with multimodal content and citations
1:  Phase 1: Query Processing
2:  $Q^{\prime}\leftarrow\text{RewriteQuery}(Q)$
3:  $\mathcal{T}\leftarrow\text{DecomposeTask}(Q^{\prime})$
4:  $\mathcal{T}_{\text{multimodal}}\leftarrow\text{ExtractMultimodalTasks}(\mathcal{T})$
5:  Phase 2: Hierarchical Retrieval
6:  for each subtask $t_{i}\in\mathcal{T}$ do
7:     $\mathcal{R}_{i}\leftarrow\text{RetrieveFromKB}(t_{i},\mathcal{K})$
8:     $\mathcal{R}_{i}^{\text{text}}\leftarrow\text{FilterTextContent}(\mathcal{R}_{i})$
9:     $\mathcal{R}_{i}^{\text{visual}}\leftarrow\text{FilterVisualContent}(\mathcal{R}_{i})$
10:  end for
11:  Phase 3: Content Organization
12:  $\mathcal{S}\leftarrow\text{GenerateSectionTitles}(\mathcal{T})$
13:  $\mathcal{C}\leftarrow\text{ClusterBySections}(\mathcal{R}^{\text{text}},\mathcal{R}^{\text{visual}},\mathcal{S})$
14:  Phase 4: Section-by-Section Writing
15:  $P\leftarrow\emptyset$, $\mathcal{H}\leftarrow\emptyset$
16:  for each section $s_{j}\in\mathcal{S}$ do
17:     $\mathcal{C}_{j}\leftarrow\text{GetSectionContent}(\mathcal{C},s_{j})$
18:     $d_{j}\leftarrow\text{CreateDraft}(s_{j},\mathcal{C}_{j},\mathcal{H})$
19:     $d_{j}^{\prime}\leftarrow\text{RefineDraft}(d_{j},\mathcal{C}_{j})$
20:     $P\leftarrow P\cup\{d_{j}^{\prime}\}$
21:     $\mathcal{H}\leftarrow\mathcal{H}\cup\text{SummarizeSection}(d_{j}^{\prime})$
22:  end for
23:  Phase 5: Multimodal Integration
24:  $\mathcal{M}\leftarrow\text{CollectMultimodalContent}(\mathcal{R}^{\text{visual}})$
25:  for each visual element $m_{k}\in\mathcal{M}$ do
26:     $\text{score}_{k,\cdot}\leftarrow\text{ComputeRelevance}(m_{k},P)$
27:     $\text{pos}_{k}\leftarrow\text{OptimizePlacement}(\text{score}_{k,\cdot},P)$
28:     $P\leftarrow\text{InsertVisual}(P,m_{k},\text{pos}_{k})$
29:  end for
30:  Phase 6: Citation Generation
31:  for each claim $c_{l}$ in $P$ do
32:     $\text{source}_{l}\leftarrow\text{FindSource}(c_{l},\mathcal{R})$
33:     $\text{citation}_{l}\leftarrow\text{GenerateCitation}(\text{source}_{l})$
34:     $P\leftarrow\text{AddCitation}(P,c_{l},\text{citation}_{l})$
35:  end for
36:  Phase 7: Final Assembly
37:  $P\leftarrow\text{AddImagePaths}(P)$
38:  $P\leftarrow\text{ValidateCitations}(P)$
39:  $P\leftarrow\text{FormatDocument}(P)$
40:  return  $P$
Online Generation Stage
During the online stage, we aim to generate a long-form, multimodal, factual grounded article based on user query $Q$ and the knowledge base $\mathcal{B}$ generated in offline stage. The overall algorithm is introduced in Algorithm 1.
Firstly, the initial user query $Q$ is rewritten and expanded to better reflect the user’s intention, which mitigates ambiguity such as using abbreviations or general questions. After rewriting the user’s query, we decompose the task according to Fact, Data and Point. The fact contains new information or restricted information. For example, when talking about "What is the total trade volume of the world in 2024?", we need to specify the countries/regions that account for, which requires the agent to generate a subtask on which countries should be included. The data directly quantifies the performance, trend or severity, which connects the fact and provides a strong foundation for the point. The task decomposition module carefully asks questions to retrieve exact numerical values to improve reliability. Besides text data, there are data illustrated in tables, charts or figures. So, we also require the task decomposition module to design strategies for searching multimodal contents that are related to the topic. The Point are conclusion drawn by editors or authors, these conclusions are considered as a viewpoint as a reference.
Next, we perform retrieval according to decomposed sub-tasks or sub-queries based on the knowledge base. To enable unified multimodal retrieval, we utilize GME Zhang et al. (2025), a multimodal embedding model that supports images and texts as the inputs. We use cosine similarity as the metric to select top-$k$ relevant documents. The relevant documents are then clustered based on pre-generated section titles, which reduces information seeking cost when generating section contents.
Now, we perform the writing process for each section. We ask the agent to write section by section according to relevant documents, focusing on combining the facts, data and points and writing in a general-to-specific pattern, such that audiences can focus on different parts, achieving customization capability. To improve fluency, we require the agent to make a draft instead of writing directly. We also summarize the written part as a compressed context to prevent repeated generated content.
Finally, we put everything together. We integrate multimodal contents by compute the relevance among different modalities. Then, we add citations for each section, and these citations are fine-grained such that we can trace to their original paragraphs. We detail these two parts in the following two subsections.
3.3 Interleaved Image-Text Generation
One challenge in generating multimodal articles lies in how to find proper places to insert multimodal contents. For text contents, we can ask the LLM to quote it based on memory mechanism or the instruction following capability. However, for multimodal contents such as images, tables, and charts, determining their optimal placement within the text requires sophisticated reasoning about content relevance and contextual flow.
To address this challenge, we propose a two-stage approach: content relevance scoring and contextual placement optimization. First, we compute relevance scores between each multimodal element and every paragraph in the generated text using the same GME embedding model. This creates a relevance matrix that captures semantic similarity between visual and textual content.
Second, we employ a placement optimization algorithm that considers both relevance scores and document flow constraints. The algorithm ensures that images are placed near their most relevant textual descriptions while maintaining the logical structure of the document. For tables and charts, we implement additional constraints to ensure they appear before or after the paragraphs that reference their data. The algorithm is shown in Algorithm 3.
3.4 Grounded Citation
A critical aspect of DeepWriter is its ability to provide precise, traceable citations that enable readers to verify claims and access source materials. Unlike traditional citation systems that reference entire documents, DeepWriter implements fine-grained citation that can point to specific pages, or even the paragraphs.
Our citation system operates at three levels of granularity: document-level, paragraph-level, and sentence-level. Each citation includes metadata such as source file name, page number, and chunk bounding box, allowing for precise source verification. For multimodal content, citations also include bounding box coordinates for images and tables.
The citation generation process is integrated into the writing pipeline, automatically creating citations as content is generated. This ensures that every factual claim, numerical value, or visual element is properly attributed to its source in the corpus. The system also maintains a citation consistency check to ensure that similar claims across different sections reference the same sources when appropriate. The algorithm is shown in Algorithm 4.
4 Experiments
4.1 System Configuration
We use MinerU as PDF processing toolkit to process a bundle of PDFs. The preprocess data are stored in a Milvus database, besides the processed information, we also add the embedding generated via GME (gme-Qwen2-VL-2B-Instruct) Zhang et al. (2025) and a descriptive caption for images and tables via a advanced vision language model Qwen2.5-VL 7B Bai et al. (2025). We use Qwen2-7B Yang et al. (2024) as the LLM performing writing tasks.
4.2 Task & Benchmark
We evaluate the performance of DeepWriter on WTR dataset [^2] See https://www.wto.org/english/res_e/reser_e/wtr_e.htm for details. WTR (World Trade Report) contains the report produced by World Trade Organization, it covers the statistic details from 2001 to 2024. We download the 23 annual reports as the corpus $\mathcal{K}$ and ask DeepWriter to write a long article for a specific topic $Q$.
For evaluation, following Shao et al. (2024); Jiang et al. (2024); Xi et al. (2025), we adopt Prometheus2-7B Kim et al. (2024) as the judge model to evaluate the quality of the generate article from several dimensions including Interest Level; Coherence and Organization, Relevance and Focus; Coverage. The evaluation standard is given in Table 2.
4.3 Models
We compare DeepWriter against the following models:
  1. 1.
    Long-Context MLLMs: We use state-of-the-art LLMs with long context windows to assess their zero-shot long-form writing capabilities. We use Qwen-Plus as the LLM to be compared in this paper, of which context window is 131,072 with maximum output of 8,192 tokens [^3] See the official website  https://www.alibabacloud.com/help/en/model-studio/what-is-qwen-llm for more details.. We also [TODO]
  2. 2.
    Naive RAG: A standard RAG implementation to evaluate its performance on long-document generation. In this setting, we directly retrieve the relevant information based on original query, then ask the LLM to write an article based on retrieved information.
  3. 3.
    Specialized Writing Models: We also compare the performance of DeepWriter against other long-form article generation systems, including STORM, CO-STORM.
The comparisons for different systems is shown in Table 1.
Table 1: Configuration for different systems
System Qwen-Plus Qwen-Plus (w titles) STORM CO-STORM DeepWriter
Model Qwen-Plus Qwen-Plus GPT-4o GPT-4o Qwen2-7B
Source Internal Internal search search offline KB
Generation single turn single turn multi-turn multi-turn multi-turn
web search no no yes yes no
5 Results
5.1 Performance on WTR dataset
Figure 3: Performance of DeepWriter on WTR dataset against Qwen-Plus, STORM and CO-STORM.
The performance on WTR dataset is given in Fig 3. It can be seen that DeepWriter achieves comparable performance with writing assistants that adopt GPT-4o as foundation models. This shows the potential of utilizing compact models to perform long article writing tasks. We also find that when adding the titles generated by DeepWriter to Qwen-Plus, the coherence improved, which illustrates the effectiveness of task decomposition and planning.
However, there are also limitations when adopting compact models. For example, except for Interest Level and Coherence and Organization, DeepWriter performs worse in the other three dimensions, which reveals the gap between small and large models.
5.2 Case Study
In this section, we conduct case studies on success and failure cases to provide deeper insights into DeepWriter’s capabilities and limitations.
Success Case: Comprehensive Trade Analysis
We asked DeepWriter to generate a report on "Global Trade Patterns in Digital Services." The system successfully decomposed this complex topic into subtasks covering regional trade flows, sector-specific trends, and regulatory frameworks. The generated document included relevant charts from the WTR dataset showing trade volume trends and properly cited specific paragraphs from source documents.
The hierarchical knowledge representation enabled efficient retrieval of both high-level concepts and specific data points. The multimodal content placement algorithm correctly positioned trade flow diagrams near their corresponding textual descriptions, enhancing document readability. Fine-grained citations allowed readers to trace every claim back to specific source paragraphs.
Failure Case: Ambiguous Query Resolution
When given the query "Economic impact of recent policies," DeepWriter struggled due to the ambiguity of "recent policies" and the lack of temporal context in the corpus. The system generated a generic overview rather than focusing on specific policy impacts, resulting in lower relevance and coverage scores.
This case highlights the importance of query rewriting and the need for better temporal reasoning capabilities. The task decomposition module failed to identify the need for temporal constraints, leading to overly broad retrieval that diluted the document’s focus.
Analysis of Limitations
Our case studies reveal several areas for improvement. First, the system requires more sophisticated temporal reasoning to handle queries involving time-sensitive information. Second, the multimodal placement algorithm sometimes struggles with complex visual elements that span multiple topics. Finally, the citation system, while precise, can become verbose for documents with many small claims.
These insights inform our future work directions, particularly the development of more advanced query understanding and temporal reasoning capabilities.
6 Conclusion
In this paper, we introduced DeepWriter, a multimodal writing assistant designed to generate high-quality, long-form documents from a curated, offline knowledge base. We demonstrated that by integrating a hierarchical knowledge structure with a robust pipeline—encompassing task decomposition, multimodal retrieval, and fine-grained citation—DeepWriter can produce factually grounded and coherent articles in specialized domains. Our experiments show that our system achieves competitive performance, particularly in factual accuracy, against strong baselines that rely on larger models or online search, validating the effectiveness of a controlled, offline approach.
Limitations & Future Work
While DeepWriter demonstrates strong performance using offline corpora, several limitations remain that provide opportunities for future research.
Knowledge Limitations
DeepWriter’s knowledge is inherently limited by the content of the curated corpus. This restriction ensures reliability but may miss recent developments or emerging trends. Future work will explore hybrid approaches that combine the reliability of offline knowledge bases with the timeliness of online information sources, implementing robust fact-checking mechanisms to maintain quality standards.
Temporal Reasoning
As demonstrated in our case studies, the current system lacks sophisticated temporal reasoning capabilities. Future versions will incorporate temporal understanding to handle queries involving time-sensitive information, policy changes, and historical trends. This includes developing temporal annotation schemes and reasoning modules that can track changes over time within the corpus.
Multimodal Complexity
The current multimodal placement algorithm works well for simple visual elements but struggles with complex charts or diagrams that span multiple topics. Future work will develop more sophisticated visual understanding capabilities, including chart parsing, data extraction, and cross-modal reasoning to better integrate complex visual content.
Scalability and Efficiency
While our hierarchical knowledge representation improves retrieval efficiency, scaling to very large corpora remains a challenge. Future work will explore more efficient indexing strategies, distributed processing approaches, and adaptive chunking algorithms to handle corpora with millions of documents.
Personalization and Customization
The current system generates documents with a general professional tone. Future work will develop user preference modeling and style adaptation capabilities to generate documents tailored to specific audiences, writing styles, and expertise levels.
Interactive Refinement
Current DeepWriter operates in a single-pass generation mode. Future work will explore interactive refinement capabilities, allowing users to provide feedback and guide the generation process iteratively, similar to human writing workflows.
These limitations represent natural extensions of our current work and provide clear research directions for advancing the field of AI-assisted document generation toward more sophisticated, reliable, and user-friendly systems.
Appendix A Prompt of DeepWriter
Query Rewriting Prompt Rewrite the following user query in a way that makes it more effective and precise. The new query should be more specific, focused, and clear, using terminology that is likely to lead to a more accurate understanding of the user’s intent. Ensure that the rewritten query captures the essence of the user’s question while improving its clarity and precision. Query: {query} Your rewritten query:
Task Decomposition Prompt You are an expert research assistant. I need to retrieve information from a database to answer the following query: Query: {query} Please help me decompose this query into 3-5 more specific, related sub-queries that would help gather comprehensive information to answer the main question. These sub-queries should: Cover different aspects of the main query Be specific enough for database retrieval Help gather contextual information needed for a complete answer Focus on factual information rather than opinions Format your response as a numbered list of sub-queries only. split them with a new line.
Section Title Generation Prompt You are an expert article writer tasked with generating section titles for a comprehensive report. Given the following query, generate a list of section titles that would be appropriate for a comprehensive report. Query: {query} Instructions: 1. The section titles should follow the human-like structure of a report. 2. The content of the section should be related to the query. 3. The section titles should from general to specific. Like: Background, Analysis, Viewpoints. 4. split the section titles by new line such that each line contains exactly one section title. Example: Background Analysis Viewpoints Your section titles:
Section Draft Prompt You are an expert research writer tasked with creating a section draft for a section of a comprehensive report. Query: {query} Section Title: {section_title} Relevant Documents: {relevant_docs} Instructions: 1. Analyze the query and section title and figure out what should be included in this section. 2. Create a rough draft for writing this section that covers the information revealed by relevant documents. 3. Be simple and concise. 4. DO NOT add references to the draft. 5. Try to avoid using bullets and subsections, just synthesize the information in a natural way. Your draft should provide a high-level perspective on how to approach writing this section effectively. Focus on organization and content strategy rather than specific wording. Provide your draft below:
Document Clustering Prompt You are an expert document classifier. Your task is to classify the given document into the most appropriate section based on its content and relevance to the query. Query: {query} Document: {doc} Available sections: {sections} Instructions: 1. Carefully analyze the document content in relation to the query 2. Consider how the information would fit into a structured report addressing the query 3. Choose EXACTLY ONE section from the available sections where this document would be most appropriate 4. Return ONLY the name of the chosen section, with no additional text or explanation Your classification (return only the section name):
Section Content Generation Prompt You are an expert research writer tasked with generating high-quality content for a specific section of a comprehensive report. Query: {query} Section Title: {section_title} Section Draft: {section_draft} Relevant Documents: {relevant_docs} Content Already Written in Previous Sections: {already_written} Instructions: 1. Generate detailed, well-structured content for the "section_title" section that directly addresses the query 2. Incorporate information from the relevant documents, synthesizing and analyzing the data 3. Ensure continuity with content already written in previous sections 4. Use an academic, professional tone appropriate for a research report 5. Be thorough but concise, focusing on information that is most relevant to the query 6. Avoid repetition of content already covered in previous sections 7. Do not include title in any level just write the content Your content should: Present factual information directly derived from the relevant documents Synthesize and organize information from multiple sources Maintain neutrality when presenting evidence and data
Summarization Prompt You are an expert summarizer. Your task is to create a concise and accurate summary of the following content in relation to a specific query. The summary should: 1. Capture the main points and key information relevant to the query 2. Highlight the relationship between the content and the query, if there is no relationship, return "None" 3. Maintain the original meaning and intent 4. Be clear and coherent 5. Be no more than 30 percent of the original length Query: {query} Content to summarize: {doc} Provide your summary below, focusing on aspects that address the query:
Appendix B Key Subroutines
Algorithm 2 shows the algorithm on task decomposition. Algorithm 3 demonstrates the multimodal placement optimization. Algorithm 4 reveals how fine-grained citation generation works.
Algorithm 2 Task Decomposition
0:  Query $Q^{\prime}$
0:  Subtasks $\mathcal{T}=\{\mathcal{T}_{\text{fact}},\mathcal{T}_{\text{data}},\mathcal{T}_{\text{point}}\}$
1:  $\mathcal{T}_{\text{fact}}\leftarrow\emptyset$
2:  $\mathcal{T}_{\text{data}}\leftarrow\emptyset$
3:  $\mathcal{T}_{\text{point}}\leftarrow\emptyset$
4:  Extract Facts
5:  for each entity $e$ in $Q^{\prime}$ do
6:     if $\text{IsAmbiguous}(e)$ then
7:        $\mathcal{T}_{\text{fact}}\leftarrow\mathcal{T}_{\text{fact}}\cup\{\text{Clarify}(e)\}$
8:     end if
9:  end for
10:  Extract Data Requirements
11:  for each quantitative term $q$ in $Q^{\prime}$ do
12:     $\mathcal{T}_{\text{data}}\leftarrow\mathcal{T}_{\text{data}}\cup\{\text{Quantify}(q)\}$
13:  end for
14:  Extract Points
15:  $\mathcal{T}_{\text{point}}\leftarrow\text{IdentifyKeyArguments}(Q^{\prime})$
16:  return  $\mathcal{T}$
Algorithm 3 Multimodal Placement Optimization
0:  Visual elements $\mathcal{M}$, Document $P$
0:  Optimized placement positions
1:  $\text{Sim}\leftarrow\emptyset$
2:  for each visual $m_{k}\in\mathcal{M}$ do
3:     for each paragraph $p_{i}\in P$ do
4:        $\text{score}_{k,i}\leftarrow\text{GME\_Similarity}(m_{k},p_{i})$
5:        $\text{Sim}[k][i]\leftarrow\text{score}_{k,i}$
6:     end for
7:  end for
8:  Placement Optimization
9:  for each visual $m_{k}\in\mathcal{M}$ do
[…]
12:     if constraints.satisfied then
13:        $_{k}\leftarrow\text{best\_pos}$_pos
14:     else
15:        $\text{pos}_{k}\leftarrow\text{FindAlternativePosition}(m_{k},\text{constraints})$
16:     end if
17:  end for
18:  return  $\{\text{pos}_{k}\}_{k=1}^{|\mathcal{M}|}$
Algorithm 4 Fine-grained Citation Generation
0:  Claim $c$, Retrieved content $\mathcal{R}$
0:  Citation with precise source location
[…]
3:  for each retrieved item $r\in\mathcal{R}$ do
[…]
8:     end if
9:  end for
[…]
16:  else
[…]
18:  end if
19:  return  citation
Appendix C LLM Evaluation
See Table 2 for details.
Criteria Description Interest Level: How engaging and thought-provoking is the article?
Score 1 Description Not engaging at all; no attempt to capture the reader’s attention.
Score 2 Description Fairly engaging with a basic narrative but lacking depth.
Score 3 Description Moderately engaging with several interesting points.
Score 4 Description Quite engaging with a well-structured narrative and noteworthy points that frequently capture and retain attention.
Score 5 Description Exceptionally engaging throughout, with a compelling narrative that consistently stimulates interest.
Criteria Description Coherence and Organization: Is the article well-organized and logically structured?
Score 1 Description Disorganized; lacks logical structure and coherence.
Score 2 Description Fairly organized; a basic structure is present but not consistently followed.
Score 3 Description Organized; a clear structure is mostly followed with some lapses in coherence.
Score 4 Description Good organization; a clear structure with minor lapses in coherence.
Score 5 Description Excellently organized; the article is logically structured with seamless transitions and a clear argument.
Criteria Description Relevance and Focus: Does the article stay on topic and maintain a clear focus?
Score 1 Description Off-topic; the content does not align with the headline or core subject.
Score 2 Description Somewhat on topic but with several digressions; the core subject is evident but not consistently adhered to.
Score 3 Description Generally on topic, despite a few unrelated details.
Score 4 Description Mostly on topic and focused; the narrative has a consistent relevance to the core subject with infrequent digressions.
Score 5 Description Exceptionally focused and entirely on topic; the article is tightly centered on the subject, with every piece of information contributing to a comprehensive understanding of the topic.
Criteria Description Broad Coverage: Does the article provide an in-depth exploration of the topic and have good coverage?
Score 1 Description Severely lacking; offers little to no coverage of the topic’s primary aspects, resulting in a very narrow perspective.
Score 2 Description Partial coverage; includes some of the topic’s main aspects but misses others, resulting in an incomplete portrayal.
Score 3 Description Acceptable breadth; covers most main aspects, though it may stray into minor unnecessary details or overlook some relevant points.
Score 4 Description Good coverage; achieves broad coverage of the topic, hitting on all major points with minimal extraneous information.
Score 5 Description Exemplary in breadth; delivers outstanding coverage, thoroughly detailing all crucial aspects of the topic without including irrelevant information.
Table 2: Scoring rubrics on a 1-5 scale for the evaluator LLM.